A disc of mass 16kg and radius 0.25m acquires an angular velocity of 126 rad/s in 8.0

seconds from rest. If the disc rotates about an axis through centre of the disc and in a

plane perpendicular to the disc, find the torque responsible for the rotation.


1
Expert's answer
2021-05-30T13:30:00-0400

"\tau=I\epsilon=\frac{1}{2}mr^2\cdot\frac{\omega}{t}=\frac{1}{2}\cdot 16\cdot0.25^2\cdot\frac{126}{8}\approx7.9\ (N\cdot m)" . Answer